### v1

# Lessons Learned from a Unifying Empirical Study of Parameter-Efficient Transfer Learning (PETL) in Visual Recognition

Abstract
Parameter-efficient transfer learning (PETL) has attracted significant attention lately, due to the increasing size of pre-trained models and the need to fine-tune them for superior downstream performance. This community-wide enthusiasm has sparked a plethora of new approaches. Nevertheless, a systematic study to understand their performance and suitable application scenarios is lacking, leaving questions like “when to apply PETL” and “which approach to use” largely unanswered, especially in visual recognition. In this paper, we conduct a unifying empirical study of representative PETL approaches in the context of Vision Transformers (ViT). We systematically tune their hyper-parameters to fairly compare their accuracy on downstream tasks. Our study not only offers a valuable user guide but also unveils several new insights. First, if tuned carefully, different PETL approaches can obtain quite similar accuracy in the low-shot benchmark VTAB-1K. This includes simple approaches like fine-tuning the bias terms that were reported inferior. Second, though with similar accuracy, we find that PETL approaches make different mistakes and high-confidence predictions, likely due to their different inductive biases. Such an inconsistency (or complementariness) opens up the opportunity for ensemble methods, and we make preliminary attempts at this. Third, going beyond the commonly used low-shot tasks, we find that PETL is also useful in many-shot regimes — it achieves comparable and sometimes better accuracy than full fine-tuning, using much fewer learnable parameters. Last but not least, we investigate PETL’s ability to preserve a pre-trained model’s robustness to distribution shifts (e.g., a CLIP backbone). Perhaps not surprisingly, PETL approaches outperform full fine-tuning alone. However, with weight-space ensembles, the fully fine-tuned model can achieve a better balance between downstream and out-of-distribution performance, suggesting a future research direction for PETL [^1] Code is available at https://github.com/OSU-MLB/PETL_Vision..
1 Introduction
Pre-training and then fine-tuning has become the standard practice to tackle visual recognition problems [7]. The community-wide enthusiasm for open-sourcing has made it possible to access large, powerful pre-trained models learned from a gigantic amount of data (e.g., ImageNet-21K [13, 77] or LAION-5B [79]). More research focus has thus been on how to fine-tune such large models [15, 96]. Among existing efforts, parameter-efficient transfer learning (PETL), a.k.a parameter-efficient fine-tuning (PEFT), has attracted increasing attention lately [25, 15]. Instead of fine-tuning the whole model (i.e., full fine-tuning) or the last fully connected layer (i.e., linear probing), PETL approaches seek to update or insert a relatively small number of parameters to the pre-trained model [94]. Doing so has several noticeable advantages. First, as named, PETL is parameter-efficient. For one downstream task (e.g., recognizing bird species or car brands), it only needs to learn and store a tiny fraction of parameters on top of the pre-trained model. Second, accuracy-wise, PETL has been shown to consistently outperform linear probing and often beat full fine-tuning, as reported on the commonly used low-shot image classification benchmark VTAB-1K [99].
(a) Accuracy gain vs. linear probing on VTAB-1K (19 tasks)
(b) In-domain (x-axis) vs. OOD (y-axis)
Figure 1: Highlights of our insights. (a) PETL methods if tuned carefully, achieve similar accuracy ($\bullet$-$\bullet$ for the range from the most to the least accurate methods) and consistently outperform linear probing ($\times$) and full fine-tuning ($\blacksquare$) on VTAB-1K. (b) Fine-tuning a CLIP ViT-B/16 with PETL on ImageNet-1K (100 samples/class) better preserves the out-of-distribution (OOD) accuracy (averaged across ImageNet-V2, ImageNet-S, ImageNet-R, and ImageNet-A) than full fine-tuning, evidenced by the $\star$ points. Interestingly, PETL is compatible with weight-space ensemble (WiSE) [92] with the pre-trained model ($\blacksquare$), but, not as effective as full fine-tuning.
To date, a plethora of PETL approaches have been proposed, bringing in inspiring ideas and promising results. Along with this come several excellent surveys that summarize existing PETL approaches [96, 94, 15]. Yet, a deeper and systematic understanding of the PETL paradigm seems still missing.
For example, with so many PETL approaches, there is a lack of unifying references for when and how to apply them. Though superior accuracy was reported on the low-shot benchmark VTAB-1K, there is not much discussion on how PETL approaches achieve it. Does it result from PETL’s ability to promote transferability or prevent over-fitting? The current evaluation also raises the question of whether PETL is useful beyond a low-shot scenario. Last but not least, besides superior accuracy, do existing PETL approaches offer different, ideally, complementary information?
Attempting to answer these questions, we conduct a unifying empirical study of representative PETL approaches in the context of Vision Transformers (ViT) [17]. This include Low-Rank Adaptation (LoRA) [34], Visual Prompt Tuning (VPT) [39], Adapter [33], and ten other approaches. We systematically tune their hyper-parameters to fairly compare their accuracy on the low-shot benchmark VTAB-1K. This includes learning rate, weight decay, and approach-specific parameters like the size of PETL parameters. Besides VTAB-1K, we examine PETL approaches on full-size downstream datasets such as CIFAR-100 [46], RESISC for remote sensing image scene classification [10], and Clevr-Distance for depth classification with synthetic data [99, 42]. We also conduct a study on ImageNet [13] and its variants with domain shifts [31, 22, 32, 76]. Our codebase will be released upon paper acceptance.
We summarize our key findings and extended analyses as follows.
Representative PETL approaches achieve similarly accuracy on VTAB-1K, if well tuned. This includes simple approaches that were reported inferior, such as fine-tuning the bias terms [98] or normalization layer coefficients [5] in the pre-trained backbone. Among all the hyper-parameters, we find the drop path rate [35] quite important. Ignoring it (i.e., setting it to 00) significantly degrades the performance. Overall, PETL approaches consistently outperform linear probing and full fine-tuning on all 19 image classification tasks (each with $1,000$ training examples) in VTAB-1K.
PETL approaches make different predictions. The above finding seems a bit disappointing: if existing PETL approaches all perform similarly in terms of accuracy, do we learn anything useful beyond a single approach? Our hunch is yes. As different PETL approaches have different inductive biases [69] — they explicitly specify different parameters to be updated or inserted for capacity control — they may learn differently from the same downstream data. Our analysis shows that different PETL approaches have different prediction errors and confidence, even though they achieve similar accuracy. This opens up the possibility of leveraging their heterogeneity for further improvement, for example, through ensemble methods [14, 104] or co-training [6, 4], and we provide preliminary studies.
PETL is also effective in many-shot regimes. We apply PETL beyond the low-shot regime and find that it is effective even with ample downstream training data — PETL can be on par or outperform full fine-tuning. This suggests that varying a fraction of the parameters of a properly chosen pre-trained backbone (e.g., pre-trained on ImageNet-21K [13, 17]) could already offer a sufficient effective capacity [100] to reach a performant hypothesis for downstream tasks.
PETL is more robust than full fine-tuning but weight-space ensemble is more effective for full fine-tuning. Besides the downstream accuracy, we also evaluate PETL’s robustness to distribution shifts, inspired by [92]. We apply PETL to fine-tune a CLIP backbone [73] with ImageNet-1K data and find that PETL preserves the backbone’s generalizability to out-of-distribution examples (e.g., from ImageNet-Sketch or ImageNet-Rendition) better than full fine-tuning. This may not be surprising. What is interesting is that the weight-space ensemble (WiSE) between the fine-tuned and pre-trained models [92] can be applied to PETL as well to further improve the robustness without sacrificing the downstream accuracy. Nevertheless, full fine-tuning with WiSE can achieve even higher accuracy in both downstream and out-of-distribution data than PETL, suggesting a further research direction in PETL.
What lead to PETL’s success? We attempt to answer this fundamental question by analyzing the results of our study. On VTAB-1K with 19 tasks, we find two cases: 1) in some tasks, full fine-tuning outperforms linear probing, suggesting the need to update the backbone; 2) in some tasks, linear probing outperforms full fine-tuning, suggesting either the backbone is good enough or updating it risks over-fitting. The superior accuracy of PETL in both cases suggests that PETL acts as an effective regularizer during low-shot training. Still using VTAB [99] but with ample training data, we find that for tasks in case 1), PETL is on par with full fine-tuning, suggesting that its regularization role does not prevent the fine-tuned model from learning sufficiently from the data. For tasks in case 2), PETL can surprisingly still outperform full fine-tuning, suggesting that it effectively transfers (or preserves) some useful pre-trained knowledge that full fine-tuning may wash away.
What do we not investigate? There are many aspects that one can ask about PETL. Our study does not consider computation-specific properties like memory usage and FLOPS.
2 Background
2.1 Large pre-trained models
Pre-trained models have become an indispensable part of modern AI development [7]. Building upon neural networks with millions if not billions of parameters and gigantic amounts of training data, these large pre-trained models have led to groundbreaking results in various downstream tasks [54, 63, 68] and shown several emerging capabilities not observed previously [44, 51, 7]. For example, in computer vision, a Vision Transformer (ViT) [17] trained with ImageNet-21K (around $14$M images) leads to consistent gains v.s. a ViT trained with ImageNet-1K (around $1.3$M images) [17]. ViTs pre-trained with millions of image-text pairs via a contrastive objective function (e.g., a CLIP-ViT model) [73, 11] show an unprecedented zero-shot capability and robustness to distribution shifts [73]. In this paper, we focus on the ImageNet-21K-ViT and use the CLIP-ViT in the robustness study.
Vision Transformer (ViT). We briefly review ViTs [17], which are adapted from the Transformer-based models [87] in NLP [91]. ViTs divide an image $\bm{I}$ into a sequence of $N$ fixed-sized patches $\{\bm{I}^{(n)}\}_{n=1}^{N}$ and treat them like NLP tokens. Each patch is first embedded into a $D$-dimensional vector ${\bm{x}}^{(n)}_{0}$ with positional encoding. The sequence of vectors is then prepended with a “CLS” vector ${\bm{x}}_{0}^{(\text{Class})}$ to generate the input $\bm{Z}_{0}=[{\bm{x}}_{0}^{(\text{Class})},{\bm{x}}^{(1)}_{0},\cdots,{\bm{x}}^{(N)}_{0}]\in\mathbb{R}^{D\times(1+N)}$ to a ViT, composed of $M$ Transformer layers. We use super-/sub-script to index token/layer. The output vector ${\bm{x}}_{M}^{(\text{Class})}$ corresponding to the “CLS” token is used as the image representation. For classification tasks, ${\bm{x}}_{M}^{(\text{Class})}$ is fed into a fully connected (FC) layer for the final prediction.
Each of the ViT’s $M$ Transformer layers consists of a multi-head self-attention (MSA) block, a multi-layer perceptron (MLP) block, two Layer Normalization (LN) blocks [3], and two residual links. The $m$-th Transformer layer can be formulated as
$\bm{Z}_{m}^{\prime}$ $=\operatorname{MSA}\left(\operatorname{LN}\left(\bm{Z}_{m-1}\right)\right)+\bm{Z}_{m-1},$ (1)
$\bm{Z}_{m}$ $=\operatorname{MLP}\left(\operatorname{LN}\left(\bm{Z}_{m}^{\prime}\right)\right)+\bm{Z}_{m}^{\prime},$ (2)
where $\bm{Z}_{m-1}=[{\bm{x}}_{m-1}^{(\text{Class})},{\bm{x}}^{(1)}_{m-1},\cdots,{\bm{x}}^{(N)}_{m-1}]\in\mathbb{R}^{D\times(1+N)}$ is the output of the preceding $(m-1)$-th Transformer layer. The MLP is applied to each column vector of $\bm{Z}_{m}^{\prime}$ independently.
Without loss of generality, let us consider an MSA block with a single head. Given a generic input $\bm{Z}\in\mathbb{R}^{D\times(1+N)}$, this block first projects it into three matrices, Query $\bm{Q}$, Key $\bm{K}$, and Value $\bm{V}$
$\bm{Q}=\bm{W}_{Q}\bm{Z},\quad\bm{K}=\bm{W}_{K}\bm{Z},\quad\bm{V}=\bm{W}_{V}\bm{Z},$ (3)
where $\bm{W}_{Q/K/V}\in\mathbb{R}^{D\times D}$ are projection matrices. The output of this block is then formulated as
$\bm{V}\times\mathsf{Softmax}(\frac{\bm{K}^{\top}\bm{Q}}{\sqrt{D}})\quad\in\mathbb{R}^{D\times(1+N)}.$ (4)
2.2 Parameter efficient transfer learning (PETL)
Fine-tuning is arguably the most common way to tailor a pre-trained model for downstream tasks. As the size of pre-trained models gets larger, copying and updating all the parameters for one downstream task becomes inefficient. PETL has thus emerged as a promising paradigm.
PETL was originally developed in NLP [26, 49, 30, 66, 82, 98, 2, 88, 56, 81, 103] and has attracted increasing attention in vision [39, 9, 40, 102, 57, 53]. Existing approaches can generally be categorized into four groups: prompt-based, adapter-based, direct selective parameter tuning, and efficient selective parameter tuning. We focus on visual recognition and compare representative PETL approaches applicable to ViTs. During fine-tuning, all approaches learn a new FC layer for prediction.
Prompt-based approaches. Prompt-based learning emerged in NLP [55, 52]. The core concept is to augment the input data with task-specific hints (prompts). Visual Prompt Tuning (VPT) [39] adapts such an idea to ViTs. Specifically, its deep version (VPT-Deep) prepends a set of soft prompts to the input tokens of each Transformer layer (i.e., $\{\bm{Z}_{m}\}_{m=0}^{M-1}$) and only optimizes the prompts during fine-tuning. Other representative works in this category include [97, 86, 24].
Adapter-based approaches. This category typically introduces additional trainable parameters (e.g., an MLP block) to the frozen pre-trained model [52]. It was initially developed for multi-domain adaptation [74, 75] and continual learning [78, 64], and was subsequently extended to the NLP and vision domains to adapt Transformer-based models [33, 97].
We consider five popular adapter-based methods. Houl. Adapter [33] is the first adapter-based PETL approach. It inserts two Adapters — a two-layer bottleneck-structured MLP with a residual link — into each Transformer layer, one after the MSA block and the other after the MLP block. Pfeif. Adapter [72] inserts the Adapter solely after the MLP block, a strategy shown effective in recent studies [34]. AdaptFormer [9] inserts the Adapter in parallel with the original MLP block in a Transformer layer, different from the sequential design of Houl. and Pfeif. Adapter. One can view it as an ensemble, summing the task-specific features (by the Adapter) and the task-agnostic features (by the original MLP) to form $\bm{Z}_{m}$ in Equation 2. ConvPass [40] introduces a convolutional-based bottleneck module (without a skip link) that explicitly encodes visual inductive biases: the 2D convolution is performed over tokens of nearby patches. The module is inserted in parallel with the MSA and/or MLP block. RepAdapter [61] introduces a linear Adapter with group-wise transformations [62] and sequentially inserts two such modules after both MSA and MLP blocks.
Direct selective parameter tuning. This category selectively updates a subset of parameters of the pre-trained model, seen as a trade-off between full fine-tuning and linear probing. We consider three approaches. BitFit [98] updates the bias terms, including those in the Q/K/V projections, the MLP blocks, the LN blocks, and the projection for patch embeddings. LayerNorm [5] updates the trainable parameters of the LN blocks in each Transformer layer. DiffFit [93] updates both the bias terms and the LN blocks and inserts learnable factors to scale the features after the MSA and the MLP blocks. Instead of updating parameters, SSF [53] linearly adapts intermediate features, motivated by feature modulation [36, 71]. For an intermediate feature $\bm{Z}\in\mathbb{R}^{D\times(N+1)}$, SSF learns a $D$-dimensional scaling vector and a $D$-dimensional additive vector broadcasting to the tokens.
Efficient selective parameter tuning. Unlike the above category which directly updates parameters, this category learns additive residuals (e.g., $\Delta\bm{W}$) to the original parameters (e.g., $\bm{W}$). By injecting a low-rank constraint to the residuals, this category effectively reduces the learnable parameters, suitable for updating parameter-heavy components like projection matrices. LoRA [34] is arguably the most well-known approach. It parameterizes the residuals by low-rank decomposition to update the Query/Value projection matrices $\bm{W}_{Q/V}\in\mathbb{R}^{D\times D}$. Concretely, to update a $\bm{W}\in\mathbb{R}^{D\times D}$ matrix, LORA learns $\bm{W}_{\text{down}}\in\mathbb{R}^{r\times D}$ and $\bm{W}_{\text{up}}\in\mathbb{R}^{D\times r}$ with $r\ll D$, and forms the additive residual by $\Delta\bm{W}=\bm{W}_{\text{up}}\bm{W}_{\text{down}}\in\mathbb{R}^{D\times D}$. Factor Tuning (FacT) [41] extends the idea of matrix decomposition into tensor decomposition. It stacks the $D\times D$ learnable matrices in all the Transformer layers into a 3D tensor and learns an additive residual parameterized by the well-established Tensor-Train (TT) [70] and Tucker (TK) [12] formulations.
More detailed descriptions of ViT and PETL methods can be founded in Appendix B.
2.3 Related work and comparison
The community-wide enthusiasm for PETL has led to multiple survey articles [96, 94, 25]. Meanwhile, several empirical, integrative, and theoretical studies were presented, mostly based on NLP tasks, attempting to provide a holistic understanding. [26, 67] provided unified views to methodologically connect PETL approaches. [8, 15, 28] and [29] empirically compared PETL approaches on NLP and vision tasks, respectively, while [20] offered a theoretical stability and generalization analysis. Accuracy-wise, [8, 15, 28] found that PETL is robust to over-fitting and quite effective in NLP tasks under low-data regimes. This is, however, not the case for vision tasks: [29] showed that representative PETL approaches like LoRA and Adapter cannot consistently outperform either full fine-tuning or linear probing. In terms of why PETL works, [20] framed PETL as sparse fine-tuning and showed that it imposes a regularization by controlling stability; [15, 29] framed PETL as (subspace) optimization; [15] further discussed the theoretical principle inspired by optimal control.
Our study strengthens and complements the above studies and offers new insights. First, we compared over ten PETL approaches, more than any of the above. We carefully tune the hyper-parameters, aiming to reveal the faithful accuracy of each approach. This is particularly important for the vision community because there have been no unifying references for PETL accuracy; simple approaches like BitFit have often been reported as quite inferior; the effectiveness of other approaches was reported quite discrepant from the study in NLP. Second, we go beyond a competition perspective to investigate a complementary perspective of PETL approaches. We show that different PETL approaches offer effective base learners for model ensembles. Third, we go beyond downstream accuracy to investigate PETL’s effectiveness in maintaining out-of-distribution robustness. Fourth, we systematically analyze the results from low-shot and many-shot regimes and identify two distinct patterns among PETL, full fine-tuning, and linear probing, extending the understanding of PETL.
3 Results on Low-Shot Regimes
Pre-trained models are meant to ease downstream applications. One representative scenario is low-shot learning: supervised fine-tuning with a small number of examples per class. Indeed, low-shot learning has been widely used to evaluate PETL performance.
Dataset. VTAB-1K [99] consists of 19 image classification tasks from three groups. The Natural group comprises natural images captured with standard cameras. The Specialized group contains images captured by specialist equipment for remote sensing and medical purposes. The Structured group evaluates the scene structure comprehension, such as object counting and 3D depth estimation. Following [99], we perform an 80/20 split on the 1000 training images in each task for hyperparameter searching. The reported result (top-1 classification accuracy) is obtained by training on the 1000 training images and evaluating on the original test set. We provide more detail in the Appendix.
Methods. We consider linear probing, full fine-tuning, and 13 PETL methods including 1 prompt-based [39], 5 adapter-based [33, 72, 9, 40, 61], 4 Direct selective [98, 5, 93, 53], and 3 Efficient selective [34, 41]. Please refer to subsection 2.2 for details.
Setup. We employ the ViT-B/16 model [17] pre-trained on ImageNet-21K [13] as the backbone. The prediction head is randomly initialized for each dataset. Images are resized to $224\times 224$. We perform no data augmentation and normalization during training, following [99]. (The exception is CIFAR-100 [46], on which normalization is necessary.) We systematically tune 1) learning rate, 2) weight decay, and 3) approach-specifics like the size of PETL parameters which are often left intact. We set a cap for 3), $\leq 1.5\%$ of ViT-B/16. We also turn the drop path rate on (e.g., 0.1) or off (i.e., 0). A detailed hyper-parameter search grid and additional training details are provided in the Appendix.
Results. As shown in 1(a) and Table 1, PETL methods generally outperform both linear probing and full fine-tuning across datasets. Additionally, under fair hyper-parameter tuning, we surprisingly found that most PETL methods perform similarly as the relative standard deviations (divided by the means) in all three groups are quite low. Simple methods (e.g., Bitfit) and PETL methods originally proposed for NLP (e.g., LoRA and Adapter), which were previously reported as inferior due to un-optimized implementations and hyperparameter tuning, now demonstrate competitive performance with SOTA visual PETL methods. To understand the relative advantages of different approaches, we provide the ranking of PETL methods across different groups in Figure 2, where the element $(i,j)$ in each ranking matrix represents the frequency that method $i$ ranks $j$th in each group. Methods are ordered by their mean ranks (in brackets), and the parameters column indicates the number of trainable parameters in millions. In the natural group, methods with very few trainable parameters perform typically better. Conversely, in the specialized and structured groups, methods with more parameters generally yield better performance. We hypothesize that this performance discrepancy arises from the domain affinity between the pre-trained domain (ImageNet) and the downstream domains. The natural group, sharing a stronger affinity with ImageNet, allows simpler methods like BitFit to adjust the features effectively. In contrast, the specialized and structured groups necessitate more complex methods with more trainable parameters to bridge the domain gap.
Recipes. In low-shot regimes, when the downstream data are similar to the pre-trained data, simple methods (e.g., DiffFit) with decent accuracy and much fewer parameters are preferred. When there is a large domain gap, complex methods with higher accuracy (but more parameters) become competitive. Learning with low-shot data is prone to over-fitting. We find that if the drop path rate — which stochastically drops a transformer block per sample [35] — is set not as default (i.e., nonzero), all the methods can benefit from such a regularization, as shown in Figure 10 in the Appendix.
Table 1: Results on VTAB-1K (19 tasks from 3 groups). We include VPT-Shallow [39] just for reference. Based on the accuracy among PETL, linear probing, and full fine-tuning, we find two task groups (purple and orange).
Natural Specialized Structured
Method
CIFAR-100
Caltech101
DTD
Flowers102
Pets
SVHN
Sun397
Mean
Camelyon
EuroSAT
Resisc45
Retinopathy
Mean
Clevr-Count
Clevr-Dist
DMLab
KITTI-Dist
dSpr-Loc
dSpr-Ori
sNORB-Azim
sNORB-Elev
Mean
Overall Mean
Tunable Params
Linear 78.1 86.6 65.7 98.9 89.3 41.5 53.2 72.5 83.1 90.0 74.9 74.6 80.6 37.5 35.1 36.5 64.6 16.2 29.4 17.3 23.7 32.5 61.9 0
Full 62.4 89.9 61.9 97.4 85.8 88.9 36.8 76.7 81.6 88.1 81.6 73.6 81.2 56.2 60.9 48.2 77.9 68.5 46.6 31.0 28.3 52.2 70.0 85.8
VPT-Shallow 80.2 88.7 67.9 99.1 89.6 77.0 54.2 79.4 81.8 90.3 77.2 74.4 80.9 42.2 52.4 38 66.5 52.4 43.1 15.2 23.2 41.6 67.3 0.07
VPT-Deep 84.8 91.5 69.4 99.1 91.0 85.6 54.7 81.8 86.4 94.9 84.2 73.9 84.9 79.3 62.4 48.5 77.9 80.3 56.4 33.2 43.8 60.2 75.6 0.43
BitFit 86.5 90.5 70.3 98.9 91.0 91.2 54.2 82.6 86.7 95.0 85.3 75.5 85.6 77.2 63.2 51.2 79.2 78.6 53.9 30.1 34.7 58.5 75.6 0.1
DiffFit 86.3 90.2 71.2 99.2 91.7 91.2 56.1 83.2 85.8 94.1 80.9 75.2 84.0 80.1 63.4 50.9 81.0 77.8 52.8 30.7 35.5 59.0 75.4 0.14
LayerNorm 86.0 89.7 72.2 99.1 91.4 90.0 56.1 83.0 84.7 93.8 83.0 75.2 84.2 77.5 62.2 49.9 78.1 78.0 52.1 24.3 34.4 57.1 74.7 0.04
SSF 86.6 89.8 68.8 99.1 91.4 91.2 56.5 82.8 86.1 94.5 83.2 74.8 84.7 80.1 63.6 53.0 81.4 85.6 52.1 31.9 37.2 60.6 76.0 0.21
Pfeif. Adapter 86.3 91.5 72.1 99.2 91.4 88.5 55.7 83.0 86.2 95.5 85.3 76.2 85.8 83.1 65.2 51.4 80.2 83.3 56.6 33.8 41.1 61.8 76.9 0.67
Houl. Adapter 84.3 92.1 72.3 98 91.7 90.0 55.4 83.2 88.7 95.3 86.5 75.2 86.4 82.9 63.6 53.8 79.6 84.4 54.3 34.2 44.3 62.1 77.2 0.77
AdaptFormer 85.8 91.8 70.5 99.2 91.8 89.4 56.7 83.2 86.8 95.0 86.5 76.3 86.2 82.9 64.1 52.8 80.0 84.7 53.0 33.0 41.4 61.5 76.9 0.46
RepAdapter 86.0 92.5 69.1 99.1 90.9 90.9 55.4 82.9 86.9 95.3 86.0 75.4 85.9 82.5 63.5 51.4 80.2 85.4 52.1 35.7 41.7 61.6 76.8 0.53
Convpass 85.0 92.1 72.0 99.3 91.3 90.8 55.9 83.5 87.7 95.8 85.9 75.9 86.3 82.3 65.2 53.8 78.1 86.5 55.3 38.6 45.1 63.1 77.6 0.49
LoRA 85.7 92.6 69.8 99.1 90.5 88.5 55.5 82.6 87.5 94.9 85.9 75.7 86.0 82.9 63.9 51.8 79.9 86.6 47.2 33.4 42.5 61.0 76.5 0.55
FacT_TT 85.8 91.8 71.5 99.3 91.1 90.8 55.9 83.4 87.7 94.9 85.0 75.6 85.8 83.0 64.0 49.0 79.3 85.8 53.1 32.8 43.7 61.3 76.8 0.13
FacT_TK 86.2 92.5 71.8 99.1 90.1 91.2 56.2 83.4 85.8 95.5 86.0 75.7 85.8 82.7 65.1 51.5 78.9 86.7 53.1 27.8 40.8 60.8 76.6 0.23
Relative Std Dev 0.81 1.13 1.78 0.34 0.54 1.82 1.24 0.54 1.20 0.59 1.95 0.83 0.94 2.67 1.50 3.22 1.37 4.11 4.46 11.02 9.30 2.70 1.09 -
Refer to caption Refer to caption Refer to caption
Figure 2: Ranking 15 methods (14 PETL including VPT-Shallow [39], and linear probing) in three groups of VTAB-1K tasks. Element $(i,j)$ is the number of times method $i$ ranks $j$th in each group. Methods are ordered (top-down) by mean ranks (in brackets). The parameters column shows the # of trainable parameters in millions.
4 PETL Approaches Offer Complementary Information
The previous section shows all PETL methods achieve similar accuracy in various domains. Does it mean all PETL methods learn similar knowledge from the downstream data? As different PETL approaches have different inductive biases — they explicitly select specific parameters to update or insert different modules in various locations — we hypothesize that they may learn differently from the same downstream data.
We start by analyzing their prediction similarity on the same dataset in VTAB-1K. It is expected that their predictions are similar for datasets with very high accuracy, such as Flowers102 (avg 99.1%) and Caltech101 (avg 91.4%). Beyond them, we find that most PETL methods show diverse predictions in other datasets in VTAB-1K. Figure 5 shows the prediction similarities between 14 PETL methods in DTD, Retinopathy, and DMLab, which are from natural, specialized, and structured groups, respectively. In DTD and Retinopathy, most methods show around 20% different predictions while in DMLab, this number goes up to around 35%, even though they achieve similar accuracy.
Such diverse predictions across methods open up the possibility of leveraging their heterogeneity for further improvement. The most straightforward approach is ensemble [23, 48], e.g., average logits over methods. Figure 3 demonstrates the ensemble performance gain over all the PETL methods in each dataset, where we use the worst PETL method as the baseline. Due to the diverse predictions across methods, the ensemble can provide consistent gain.
Additionally, we analyze if PETL methods make similar correct predictions for high-confidence samples and similar mistakes for low-confidence samples. Figure 5 shows the correct prediction overlap for the 5K most confident samples (per method) and the wrong prediction overlap for the 5K least confident samples (per method). For demonstration purposes, we select one method from each PETL category (LoRA, Adapter, SSF) and they are fine-tuned on CIFAR-100 in VTAB-1K. Since they make different predictions in both high and low-confidence regimes, this paves the way for new possibilities of using different PETL methods to generate diverse pseudo-labels for semi-supervised learning [95, 21, 90], domain adaptation [19, 84, 18], continual learning [80, 59, 65] and transfre learning [105, 85, 58].
Figure 3: Ensemble (avg logits) provides consistent gain on most datasets thanks to the diverse prediction.
Refer to caption Refer to caption Refer to caption
Figure 4: Prediction similarity analysis: element $(i,j)$ shows the percentage of samples that method $i$ and method $j$ have the same predictions. Although different methods achieve similar overall accuracy, they have quite diverse predictions.
Figure 5: LoRA, Adapter and SSF fine-tuned on CIFAR100 in VTAB-1K. (a): the correct prediction overlap for the 5K most confident samples. (b): the wrong prediction overlap for the 5K least confident samples.
5 Results on Many-Shot Regimes
Recent works in NLP [8] have indicated that PETL methods may not perform as competitively as full fine-tuning when data is abundant. We thus aim to investigate PETL’s performance in many-shot regimes by addressing the following questions: (1) Should we use PETL or full fine-tuning when data is sufficient? (2) How should we adjust the number of trainable parameters for PETL methods in many-shot regimes?
Dataset. We select one representative dataset from each of the natural, specialized, and structured groups in VTAB: (1) CIFAR-100 [46], a natural image dataset comprising 50K training images across 100 classes; (2) RESISC [10], a remote sensing dataset for scene classification with 25.2K training samples across 45 classes; and (3) Clevr-Distance [99, 42], a synthetic image dataset for predicting the depth of the closest object from the camera with 6 depth classes and 70K samples. The reported results are obtained by training on the full training set and evaluating on the original test set.
Setup. The model setup follows the VTAB-1K experiment. We apply horizontal flipping for CIFAR100, horizontal and vertical flipping for Resisc, and no augmentation for Clevr.
Results. In many-shot regimes, with sufficient downstream data, full fine-tuning may catch up and eventually outperform PETL methods. However, from Figure 6, we, found that even in many-shot regimes, PETL can achieve quite comparable results with full fine-tuning, even just using $2\%$ of fine-tuning parameters. (The performance gain, however, quickly diminishes and plateaus after 5% of tunable parameters.) By comparing the results on the domain-close CIFAR-100 and domain-different RESISC and Clevr, we have some further observations. On the one hand, downstream tasks with larger domain gaps suggest the need to update, perhaps many, parameters to obtain high accuracy. With sufficient downstream data, full fine-tuning is less prone to over-fitting and indeed attains a high accuracy. But interestingly, PETL methods, with only $2\sim 5\%$ of tunable parameters, achieve similar accuracy, suggesting that its design principle does offer sufficient effective capacity for the model to learn [100]. On the other hand, downstream tasks with smaller domain gaps suggest that the pre-trained model had learned sufficient knowledge about them; fully fine-tuning it thus risks washing such knowledge away. In fact, we found that PETL notably outperforms full fine-tuning on CIFAR-100, suggesting it as a more robust transfer learning algorithm for downstream tasks.
Recipes. In many-shot regimes, PETL methods with sufficient parameters ($2\sim 5\%$) appear more favorable than full fine-tuning and linear probing. On the one hand, they achieve comparable and even better accuracy than full fine-tuning. On the other hand, the tunable parameters remain manageable. The parameter efficiency of PETL also often implies less training GPU memory usage and training time, making PETL methods a favorable alternative in many-shot regimes. For a downstream domain that is close to the pre-training domain, PETL shows much pronounced transferability. For a downstream domain that is quite different, the limited tunable parameters (controversially, $2\sim 5\%$ already amounts to a few million) already allow the model to learn sufficiently.
Refer to caption Refer to caption Refer to caption
Figure 6: PETL accuracy in many-shot regimes, with different parameter sizes on three datasets from different domains. Even a few trainable parameters allow the model to have sufficient capacity to learn from full data.
6 Why Do PETL Methods Work?
Within 19 VTAB-1K tasks, we see two cases: (1) Full fine-tuning outperforms linear probing. Since linear probing reflects the pre-trained feature quality for downstream tasks, case (1) suggests the necessity to update the backbone to address the gap between pre-trained and downstream domains. (2) Linear probing surpasses full fine-tuning, suggesting the pre-trained features are good enough (at least in a low-shot scenario). Recklessly updating them with limited data may risk over-fitting.
Figure 7 (a-b) summarizes the low-shot accuracy comparison based on the categorization above. Specifically, each line corresponds to one task. Linear probing, PETL, and fine-tuning are located in order, from left to right, to reflect their tunable parameter sizes. PETL’s superior performance in both cases showcases its capacity to learn and its regularization role to prevent over-fitting.
We also draw the many-shot accuracy in Figure 7 (c-d) based on the same categorization: RESISC and Clevr in case (1), and CIFAR-100 in case (2). In the many-shot setting, full fine-tuning consistently outperforms linear probing, which seems to suggest no more risk of over-fitting. However, on CIFAR-100 (Figure 7 (d)), we again see a noticeable gap between PETL and full fine-tuning, just like in Figure 7 (b). Such a concave shape reminds us of the long-standing under-fitting-over-fitting curve, suggesting that even with sufficient downstream data, full fine-tuning still risks over-fitting.
Taking into account PETL’s comparable performance to full fine-tuning on RESISC and Clevr with large domain gaps, we conclude — PETL succeeds as a high-capacity learner equipped with an effective regularizer; the two roles trade-offs well such that PETL can excel in both low-affinity and high-affinity domains under both low-shot and many-shot settings.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) (b) (c) (d)
Figure 7: (a): VTAB-1K tasks in case 1, PETL $>$ full $>$ linear . (b) VTAB-1K tasks in case 2, PETL $>$ linear $>$ full. (c) RESISC & Clevr in case 1 with sufficient data, PETL $\approx$ full $>$ linear. (d) CIFAR in case 2 with sufficient data, PETL $>$ full $>$ linear. Within each figure, left for linear, middle for PETL, and right for full.
7 Empirical Results on Robustness to Distribution Shifts
Large pre-trained models such as CLIP [73] and ALIGN [38] have demonstrated unprecedented accuracy across a range of data distributions when performing zero-shot inference. However, recent studies [92, 73] have shown that fine-tuning on downstream data, while significantly boosting performance on the target distribution, often compromises the model’s robustness to distribution shifts. Given that PETL only updates a limited number of parameters in the model, we investigate whether PETL can offer a more robust alternative to full fine-tuning for pre-trained models.
Dataset. We use 100-shot ImageNet-1K as our target distribution, with each class containing 100 images. Following [92], we consider 4 natural distribution shifts from ImageNet: ImageNet-V2 [76], a new ImageNet test set collected with the original labeling protocol; ImageNet-R [31], renditions for 200 ImageNet classes; ImageNet-S [22], sketch images for 1K ImageNet classes; ImageNet-A [32], a test set of natural images misclassified by a ResNet-50 [27] for 200 ImageNet classes.
Setup. We focus on the CLIP ViT-B/16 model, which comprises a visual encoder and a text encoder, pre-trained via contrastive learning on image-text pairs. Following [92], we add an FC layer as the prediction head with zero-initialized bias and initialize weights using the class label text embedded by the text encoder. Subsequently, we discard the text encoder and apply PETL methods to the visual encoder, fine-tuning only the PETL modules and the head.
Results. As shown in Table 2, while some PETL methods may not surpass full fine-tuning on the target distribution, they consistently demonstrate more robust performance on OOD data. This is likely because PETL updates only a small fraction of the parameters, thus preserving the robust features of the foundation models. Given the similar target distribution performance, should we blindly use PETL methods for more robustness?
Weight-space ensembles (WiSE) for PETL. WiSE [92], which linearly interpolates the full fine-tuned and original models, is a popular fine-tuning approach to enhance robustness. We explore whether WiSE can enhance the robustness of PETL. To apply WiSE to PETL, we first linearly interpolate the prediction head with a mixing coefficient $\alpha$. For direct selective tuning methods (e.g.BitFit), we directly interpolate with the original model. Since most Adapter-based methods have residual connections, we can multiply the adapter modules with $\alpha$ to control their strengths. A similar approach can be applied to efficient selective methods (e.g.LoRA) as they learn additive residuals to the original parameters. As shown in 1(b), WiSE improves both the robustness and the in-distribution performance of PETL methods. Interestingly, even though full fine-tuning is generally less robust than PETL methods, applying WiSE allows it to achieve better performance in both in-distribution and OOD data, which suggests a promising research direction for robust PETL.
Full BitFit
Layer-
Norm
Houl.
Adapter
Adapt-
Former
Rep-
Adapter
Convpass LoRA FacT_TK
100-shot ImageNet 74.95 75.27 74.80 75.01 75.62 76.50 76.26 76.56 74.69
Avg. OOD Acc 42.50 55.44 55.91 56.12 56.12 56.21 54.65 55.93 56.12
Table 2: The Avg. OOD Acc shows the average performance for ImageNet-(V2, S, R, A).
8 Conclusion
We conduct a unifying empirical study of parameter-efficient fine-tuning (PETL), an emerging topic in the large model era. We have several new insights and implications, including PETL methods’ complementary expertise, suitable application regimes, and robustness to domain shifts. We expect our study to open new research directions and serve as a valuable user guide in practice.
Appendix
We provide details that are omitted from the main paper.
  • Appendix A: Experiment and dataset details
  • Appendix B: Detailed descriptions of ViT and compared methods.
  • Appendix C: Additional results not presented in main paper
  • Appendix D: Discussion about further impacts of this work
Appendix A Experiment and Dataset Details
A.1 Experiment Details
Throughout all experiments, we employ AdamW optimizer [60] with a batch size of 64 and utilize the cosine decay learning rate scheduler. We train all methods with 100 epochs for VTAB-1K and 40 epochs for many-shot experiments. For the low-shot experiments, the learning rate is tuned from [1e-3, 1e-2] and weight decay from [1e-4, 1e-3]. The method-specific hyperparameter tuning grip is shown in Table 3 with the tunable parameter range in millions. For the many-shot experiment, the learning rate is tuned from [5e-4, 1e-3] and weight decay keeps the same range of [1e-4, 1e-3]. For the robustness experiment, following [92], we set a small learning rate as 3e-5 and weight decay as 5e-3. We use a strong data augmentation following [102].
We used a workstation with eight NVIDIA RTX 6000 Ada GPUs, two AMD EPYC 9554 64-Core Processors, and 800GB of RAM.
A.2 Dataset Details
VTAB-1K
The processed VTAB-1K can be downloaded from our official code base to ensure reproducibility.
Many-shot Datasets
We perform 90/10 train-val split for CIFAR-100, RESISC and Clevr-Distance. The split details are provided in our code base for reproducibility. We apply horizontal flipping for CIFAR100, horizontal and vertical flipping for Resisc, and no augmentation for Clevr. All data are normalized by ImageNet mean and standard deviation.
Method Hyperparamters #Params (M)
VPT-Shallow Prompt Number: [5, 10, 50, 100, 200] 0.0003 $\sim$ 0.153
VPT-Deep Prompt Number: [5, 10, 50, 100] 0.046 $\sim$ 0.921
BitFit N/A 0.102
DiffFit N/A 0.140
LayerNorm N/A 0.038
SSF N/A 0.205
Pfeif. Adapter
Adapter Scale Factor: [0.01, 0.1, 1, 10]
Adapter Bottleneck: [4, 8, 16, 32]
0.082 $\sim$0.599
Houl. Adapter
Adapter Scale Factor: [0.01, 0.1, 1, 10]
Adapter Bottleneck: [4, 8, 16, 32]
0.165 $\sim$1.198
AdaptFormer
Adapter Scale Factor: [0.05, 0.1, 0.2]
Adapter Bottleneck: [4, 16, 32]
0.082 $\sim$0.599
RepAdapter
RepAdapter Scale Factor: [0.1, 0.5, 1, 5, 10]
RepAdapter Bottleneck: [8, 16, 32]
0.239 $\sim$0.903
Convpass
Convpass Scale Factor: [0.01, 0.1, 1, 10, 100]
Convpass Bottleneck: [8, 16]
Convpass Xavier Init: [True, False]
0.327 $\sim$0.664
LoRA LoRA Bottleneck: [1, 8, 16, 32] 0.036 $\sim$1.179
FacT_TT
FacT Scale Factor: [0.01, 0.1, 1, 10, 100]
FacT Bottleneck: [8, 16, 32]
0.021 $\sim$0.196
FacT_TK
FacT Bottleneck: [16, 32, 64]
FacT Scale Factor: [0.01, 0.1, 1, 10, 100]
0.030 $\sim$0.369
Table 3: Methods-specific hyperparameter tuning grip for VTAB-1K experiment.
Appendix B Background
Symbol (Abbreviation) Definition
$(H,W)$ Resolution of input images
$C$ Number of channels (input images)
$P$ Resolution of patches
$N$ Number of patches (tokens)
$N_{h}$ Number of head in each Transformer layer
$D$ Embedding dimension
$D_{h}$ Embedding dimension for single-head attention
$L_{m}$ $m$-th Transformer layer
$M$ Number of Transformer layer
$Z_{m-1}$ Input of $m$-th Transformer layer
ViT Vision Transformer
LN Layer Normalization
MSA Multi-head Self-Attention
MLP Multi-Layer Perceptron
FC Fully-connected layer
Table 4: Definitions of symbols and abbreviation used in Appendix B
B.1 Vision Transformer
Overview of ViT. Inspired by the recent success of Transformer-based models [87] in NLP [91], Vision Transformer (ViT) [17] has become widely used in computer vision. To handle 2D images, ViT divides an image $\bm{I}\in\mathbb{R}^{H\times W\times C}$ into $N$ non-overlapping patches $\{\bm{I}^{(n)}\in\mathbb{R}^{P^{2}\times C}\}_{n=1}^{N}$, where $(H,W)$ is the resolution of the input image, $C$ is the number of channels, $N=HW/P^{2}$ and $(P,P)$ is the resolution of each patch. Each patch $\bm{I}^{(n)}$ is flattened and embedded into a $D$-dimensional vector ${\bm{x}}^{(n)}_{0}$ with a trainable linear projection. Incorporating the BERT design approach [43], a “Class” token ${\bm{x}}_{0}^{(\text{Class})}$ is prepended to the sequence of embedded patches, whose output state at the last Transformer layer is utilized as the image representation. Finally, position embeddings $\mathbf{E}_{\text{pos }}\in\mathbb{R}^{{D\times(1+N)}}$ are added to preserve positional information and form the input $\mathbf{Z}_{0}\in\mathbb{R}^{D\times(1+N)}$ to the ViT, which can be formulated by:
$\bm{Z}_{0}=\left[{\bm{x}}_{0}^{(\text{Class})},{\bm{x}}^{(1)}_{0},{\bm{x}}^{(2)}_{0},\cdots,{\bm{x}}^{(N)}_{0}\right]+\mathbf{E}_{\text{pos }}$ (5)
As shown in the left part of Figure 8, a ViT typically consists of $M$ layers, denoted by $\{L_{m}\}_{m=1}^{M}$. The input $\bm{Z}_{0}$ mentioned above is fed into the first layer $L_{1}$, producing the output $\bm{Z}_{1}=L_{1}(\bm{Z}_{0})=[{\bm{x}}_{1}^{(\text{Class})},{\bm{x}}^{(1)}_{1},\cdots,{\bm{x}}^{(N)}_{1}]\in\mathbb{R}^{D\times(1+N)}$, which maintains the same size as $\bm{Z}_{0}$. Namely, $\bm{Z}_{1}$ comprises $1+N$ feature tokens, and each corresponds to the same column in $\bm{Z}_{0}$. Similarly, for $m=2,\cdots,M$, each layer $L_{m}$ takes the output of the previous layer as input and generates the output, $\bm{Z}_{m}=L_{m}(\bm{Z}_{m-1})$. Finally, the “Class” vector ${\bm{x}}_{M}^{(\text{Class})}$ in $\bm{Z}_{M}$ serves as the image feature for prediction. When dealing with classification tasks, the predicted label $\hat{y}=\mathsf{Head}({\bm{x}}_{M}^{(\text{Class})})$ is generated through a linear head (i.e., a fully-connected layer).
Details of each Transformer layer. As shown in the right part of Figure 8, each Transformer layer consists of a Multi-head Self-Attention (MSA) block, a Multi-Layer Perceptron (MLP) block, and two Layer Normalization (LN) layers [3]. Formally, a Transformer layer $L_{m}$ can be defined as
$\bm{Z}_{m}^{\prime}$ $=\operatorname{MSA}\left(\operatorname{LN}\left(\bm{Z}_{m-1}\right)\right)+\bm{Z}_{m-1}$ (6)
$\bm{Z}_{m}$ $=\operatorname{MLP}\left(\operatorname{LN}\left(\bm{Z}_{m}^{\prime}\right)\right)+\bm{Z}_{m}^{\prime}$
where $\bm{Z}_{m-1}=[{\bm{x}}_{m-1}^{(\text{Class})},{\bm{x}}^{(1)}_{m-1},\cdots,{\bm{x}}^{(N)}_{m-1}]\in\mathbb{R}^{D\times(1+N)}$ is the output of the preceding $(m-1)$-th Transformer layer. The MLP is applied to each column vector of $\bm{Z}_{m}^{\prime}$ independently.
In order to encapsulate multiple complex relationships amongst different elements in the sequence, the MSA block comprises $N_{h}$ single-head self-attention blocks. For the $i^{th}$ single-head self-attention block, an generic input $\bm{Z}$ is first projected into three matrices, namely Query $\bm{Q}^{(i)}$, Key $\bm{K}^{(i)}$, and Value $\bm{V}^{(i)}$
$\bm{Q}^{(i)}=\bm{W}^{(i)}_{Q}\bm{Z},\quad\bm{K}^{(i)}=\bm{W}^{(i)}_{K}\bm{Z},\quad\bm{V}^{(i)}=\bm{W}_{V}^{(i)}\bm{Z},$ (7)
where $\bm{W}^{(i)}_{Q/K/V}\in\mathbb{R}^{D_{h}\times D}$  [^2] For brevity, we ignore the layer index $m$ for the projection matrices $\bm{W}_{Q},\bm{W}_{K},\bm{W}_{V}$, but each layer has its own projection matrices. where $D_{h}$ is the embedding dimension for a single head self-attention block and typically set to $D/N_{h}$. The $i^{th}$ self-attention head in MSA is formulated as
$\operatorname{Attn}^{(i)}(\bm{Z})=\bm{V}^{(i)}\times\operatorname{Softmax}\left(\frac{\left.{\bm{K}^{(i)}}^{\top}\bm{Q}^{(i)}\right)}{\sqrt{D_{h}}}\right)\in\mathbb{R}^{D_{h}\times(1+N)}$ (8)
The outputs of all heads are concatenated and linearly projected by a fully connected layer ($FC_{attn}$) with weight $\bm{W}_{O}\in\mathbb{R}^{D\times(D_{h}\cdot N_{h})}$ as the output of the MSA block.
$\operatorname{MSA}(\bm{Z})=\bm{W}_{O}\left[\operatorname{Attn}^{0}(\bm{Z}),\ldots,\operatorname{Attn}^{N_{h}}(\bm{Z})\right]$ (9)
The MLP block can be defined as
$\operatorname{MLP}(\bm{Z})=GELU\left(\bm{Z}\bm{W}_{1}+b_{1}\right)\bm{W}_{2}+b_{2}$ (10)
where $\bm{W}_{1}\in\mathbb{R}^{D\times 4D}$3334 is the MLP ratio in ViT-B, $\bm{W}_{2}\in\mathbb{R}^{4D\times D}$, $b_{1}\in\mathbb{R}^{4D}$, $b_{2}\in\mathbb{R}^{D}$ are weights and biases for two FC layers ($FC_{1}$ and $FC_{2}$) respectively.
Since PETL methods often entail incorporating additional components to modify the intermediate features within or between Transformer layers, we adopt the notation $\{h_{1},\ldots,h_{10}\}$ to denote the intermediate features in the unravelled view of a Transformer layer (as depicted in Figure 8) to facilitate a clearer illustration of the PETL methods discussed in the subsequent section.
B.2 Evaluated Methods
In this section, we dive into the details of 12 State-Of-The-Art PETL approaches, categorized into three groups: Prompt-based, Adapter-based and Selective Parameter Tuning. We will describe the distinctions and tradeoffs between them. A consolidated overview of these approaches is summarized in Table 5.
Figure 8: ViT h10, h1, h6 are confusing
Table 5: PETL Methods Summary: Prompt-based and adapter-based methods incorporate additional parameters to modify features while keeping the original backbone intact. However, these added parameters introduce additional inference overhead. In contrast, selective tuning methods modify the backbone by updating selective parameters, thereby incurring no additional inference overhead.
Method What Tunable Parameters Hyper Parameters Modified Type Inference Efficient
VPT-Deep $\bm{h_{1}}=[\bm{h_{1}},\bm{P}]$ $\bm{P}\in\mathbb{R}^{l\times D}$ $l$: Number of prompts Feature
AdaptFormer $\bm{h_{9}}=\bm{h_{9}}+\operatorname{Adapter}\bm{(h_{7})}$ $\bm{W}_{\text{down }/\text{up}}\in\mathbb{R}^{r\times D/D\times r}$ in $\operatorname{Adapter}$ $s$: Scale factor in $\operatorname{Adapter}$ $r$: Bottleneck dimension Feature
Pfeif. Adapter $\bm{h_{9}}=\operatorname{Adapter}(\bm{h_{9}})$ $\bm{W}_{\text{down }/\text{up}}\in\mathbb{R}^{r\times D/D\times r}$ in $\operatorname{Adapter}$ $s$: Scale factor in $\operatorname{Adapter}$ $r$: Bottleneck dimension Feature
Houl. Adapter $\bm{h_{5}}=\operatorname{Adapter_{1}}(\bm{h_{5}})$ $\bm{h_{9}}=\operatorname{Adapter_{2}}(\bm{h_{9}})$ $\bm{W}^{1}_{\text{down }/\text{up}}\in\mathbb{R}^{r\times D/D\times r}$ in $\operatorname{Adapter_{1}}$ $\bm{W}^{2}_{\text{down }/\text{up}}\in\mathbb{R}^{r\times D/D\times r}$ in $\operatorname{Adapter_{2}}$ $s$: Scale factor in $\operatorname{Adapter}$ $r$: Bottleneck dimension Feature
Convpass $\bm{h_{5}}=\operatorname{Convpass_{1}}(\bm{h_{2}})+\bm{h_{5}}$ $\bm{h_{9}}=\operatorname{Convpass_{2}}(\bm{h_{7}})+\bm{h_{9}}$ $\bm{W}^{1}_{conv2d}\in\mathbb{R}^{r\times r\times k\times k}$ $\bm{W}^{1}_{\text{down }/\text{up}}\in\mathbb{R}^{r\times D/D\times r}$ in $\operatorname{Convpass_{1}}$ $\bm{W}^{2}_{conv2d}\in\mathbb{R}^{r\times r\times k\times k}$ $\bm{W}^{2}_{\text{down }/\text{up}}\in\mathbb{R}^{r\times D/D\times r}$ in $\operatorname{Convpass_{2}}$ $s$: Scale factor in $\operatorname{Convpass}$ $r$: Bottleneck dimension $k$: Kernel size of $conv2d$ Feature
RepAdpater $\bm{h_{2}}=\operatorname{RepAdapter_{1}}(\bm{h_{2}})$ $\bm{h_{7}}=\operatorname{RepAdapter_{2}}(\bm{h_{7}})$ $\bm{W}^{1}_{conv1d}\in\mathbb{R}^{r\times D}$ $\bm{b}^{1}\in\mathbb{R}^{r}$ in $\operatorname{RepAdapter_{1}}$ $\bm{W}^{2}_{conv1d}\in\mathbb{R}^{D\times\frac{r}{G}}$ $\bm{b}^{2}\in\mathbb{R}^{D}$ in $\operatorname{RepAdapter_{2}}$ $s$: Scale factor in $\operatorname{RepAdapter}$ $r$: Bottleneck dimension $G$: Number of groups Feature
LayerNorm $\bm{h_{2}}=\operatorname{LayerNorm_{1}}(\bm{h_{1}})$ $\bm{h_{7}}=\operatorname{LayerNorm_{2}}(\bm{h_{6}})$ $\bm{W}^{1(2)},\bm{b}^{1(2)}\in\mathbb{R}^{D}$ in $\operatorname{LayerNorm_{1(2)}}$ N/A Backbone
BitFit Fine-tune all bias terms in the network $\bm{b}^{1(2)}\in\mathbb{R}^{D}$ in $\operatorname{LayerNorm_{1(2)}}$ $\bm{b}^{Q/K/V}\in\mathbb{R}^{D}$ in $\operatorname{Q/K/V}$ $\bm{b}^{FC_{attn}}\in\mathbb{R}^{D}$ in $\operatorname{FC_{attn}}$ $\bm{b}^{1}\in\mathbb{R}^{4D},$ in $\operatorname{FC_{1}}$, $\bm{b}^{2}\in\mathbb{R}^{D}$ in $\operatorname{FC_{2}}$ N/A Backbone
DiffFit     LayerNorm + BitFit      All tunable parameters in LayerNorm & BitFit N/A Backbone
     $\bm{h_{5}}=\gamma_{1}\cdot\bm{h_{5}}$ $\bm{h_{9}}=\gamma_{2}\cdot\bm{h_{9}}$     $\gamma_{1},\gamma_{2}\in\mathbb{R}^{D}$
SSF $\bm{h_{2}}=\operatorname{SSF_{2}}(\bm{h_{2}})$, $\bm{h_{3}}=\operatorname{SSF_{3}}(\bm{h_{3}})$ $\bm{h_{5}}=\operatorname{SSF_{5}}(\bm{h_{5}})$, $\bm{h_{7}}=\operatorname{SSF_{7}}(\bm{h_{7}})$ $\bm{h_{8}}=\operatorname{SSF_{7}}(\bm{h_{8}})$, $\bm{h_{9}}=\operatorname{SSF_{9}}(\bm{h_{9}})$ $\bm{W}^{2,5,7,9}\in\mathbb{R}^{D},\bm{b}^{2,5,7,9}\in\mathbb{R}^{D}$ $\bm{W}^{3}\in\mathbb{R}^{3D},\bm{b}^{3}\in\mathbb{R}^{3D}$ $\bm{W}^{8}\in\mathbb{R}^{4D},\bm{b}^{8}\in\mathbb{R}^{4D}$ N/A Backbone
LoRA $\bm{h_{3}}=\operatorname{LoRA}(\bm{h_{2}})+\bm{h_{3}}$ $\bm{W}_{\text{down}/\text{up}}^{Q/K/V}\in\mathbb{R}^{r\times D/D\times r}$ in $\operatorname{LoRA}$ $r$: Bottleneck dimension Backbone
FacTTT(TK) $\bm{h_{3}}=\operatorname{FacT_{TT(TK)}}(\bm{h_{2}})+\bm{h_{3}}$ $\bm{h_{5}}=\operatorname{FacT_{TT(TK)}}(\bm{h_{4}})+\bm{h_{5}}$ $\bm{h_{8}}=\operatorname{FacT_{TT(TK)}}(\bm{h_{7}})+\bm{h_{8}}$ $\bm{h_{9}}=\operatorname{FacT_{TT(TK)}}(\bm{h_{8}})+\bm{h_{9}}$ $\bm{U}\in\mathbb{R}^{D\times r},\bm{V}\in\mathbb{R}^{D\times r}$, $\bm{\Sigma}\in\mathbb{R}^{12L\times r\times r}$ in $\operatorname{FacT_{TT}}$ $s$: Scale factor in $\operatorname{FacT_{TT(TK)}}$ $r$: Bottleneck dimension Backbone
$\bm{U}\in\mathbb{R}^{D\times r},\bm{V}\in\mathbb{R}^{D\times r}$, $\bm{A}\in\mathbb{R}^{12L\times r},\bm{B}\in\mathbb{R}^{r\times r\times r}$ in $\operatorname{FacT_{TK}}$
B.3 Prompt-based Methods
Prompt-based learning emerged in NLP as an effective approach to adapt pre-trained models for downstream tasks [55, 52]. The core concept involves augmenting the model input with task-specific hints (prompts), which aid the pre-trained model in addressing novel tasks with its existing knowledge. Hard prompts are human-interpretable natural language hints, encompassing task instructions, in-context examples, or supporting information. Alternatively, soft prompts are continuous vector hints that are incorporated into the input embeddings of the input layers or hidden states of other layers. Soft prompts are updated during the fine-tuning process using gradient-based methods, guided by the downstream task-specific loss functions, while the pre-trained model itself remains fixed. The splendent success of prompts in NLP has sparked a growing interest in adopting it in computer vision [97, 86] and multi-modal domains [24].
In this paper, we investigate a prominent and strong prompt-based method called Visual Prompt Tuning (VPT) [39], which represents one of the early endeavours in introducing prompts to computer vision. Specifically, VPT-Shallow adds $l$ prompts $\bm{P}_{0}\in\mathbb{R}^{l\times D}$ to the input of the first Transformer layer $\bm{Z}_{0}$ and the output $\tilde{\bm{P}_{0}}$ of $\bm{P}_{0}$ serves as the input for the next layer as depicted in Equation 11. VPT-Shallow can be perceived as the addition of learnable pixels to the original images. On the other hand, VPT-Deep inserts $l$ prompts $\{\bm{P}_{m}\in\mathbb{R}^{l\times D}\}_{m=0}^{M}$ to the input of every Transformer layer $\bm{Z}_{m}$ but their outputs are discarded at the end of the layer as illustrated in Equation 12.
$=L_{m}([\bm{P}_{0},\bm{Z}_{0}])$ (11)
$[\tilde{\bm{P}_{m}},\bm{Z}_{m}]$ $=L_{m}([\tilde{\bm{P}}_{m-1},\bm{Z}_{m-1}])\quad m=2,3,\ldots,M$
[_$[\_,\bm{Z}_{m}]=L_{m}([{\bm{P}}_{m-1},\bm{Z}_{m-1}])\quad m=1,2,3,\ldots,M\$ (12)
Throughout the adaptation process, the pre-trained model is frozen and no additional weights are introduced to the model, thereby preserving the model’s original behaviour. During the forward pass, the output $\bm{Z}_{m}$ of layer $m$ is changed because of the interaction between $\bm{Z}_{m-1}$ and $\bm{P}_{m-1}$ (or $\tilde{\bm{P}}_{m-1}$) in the MSA block. Thus, the output feature is adapted to the downstream tasks by iteratively tuning the prompts through gradient descent.
B.4 Adapter-based Methods
Adapter-based methods typically introduce additional trainable parameters into a frozen pre-trained model to facilitate learning of downstream tasks [52]. Initially developed for multi-domain adaptation [74, 75] and continual learning [78, 64], the idea of Adapters is subsequently embraced by Houlsby et al[33] in the NLP domain to adapt Transformer-based networks for downstream tasks, and it also has garnered increasing interest in the computer vision field [97]. In this comparative analysis, we concentrate on five popular Adapter-based methods, encompassing the original Adapter, along with variants focusing on adjusting the positions of Adapters [9, 72], introducing visual inductive biases [40], as well as employing re-parameterization to reduce the number of trainable parameters and inference latency [61].
Houl. Adapter
[33] inserts two lightweight bottleneck-structured modules into each Transformer layer: one after the MSA block and the other after the MLP block. As depicted in 9(a), the Adapter is composed of a down-projection layer with $\bm{W}_{\text{down }}\in\mathbb{R}^{r\times D}$, a nonlinear activation function $\sigma$, an up-projection layer with $\bm{W}_{\text{up}}\in\mathbb{R}^{D\times r}$, a scaling factor $s$ and a skip-connection. To limit the number of trainable parameters, the bottleneck dimension is much smaller than the feature dimension $r\ll D$. Formally, Houl. Adapter can be defined as:
$h_{5}=\operatorname{Adapter_{1}}(h_{5})\quad h_{9}=\operatorname{Adapter_{2}}(h_{9})$ (13)
$\operatorname{Adapter}(h)=s\cdot\bm{W}_{\text{up}}\sigma(\bm{W}_{\text{down }}h)+h$ (14)
Pfeif. Adapter
[72] is a more efficient variant that introduces the Adapter solely after the MLP block, a strategy that has demonstrated effectiveness in recent studies [34]. Pfeif. Adapter can be defined formally as $h_{9}=\operatorname{Adapter}(h_{9})$ where $\operatorname{Adapter}$ follows Equation 14.
AdaptFormer
[9] proposed to insert the Adapter in parallel with the MLP block, which differs from the sequential design of Houl. and Pfeif. Adapter. The rationale behind this parallel design lies in the belief that the domain-specific features generated by the Adapter can complement the domain-agnostic features derived from the original MLP block, leading to an improved feature ensemble [83]. Formally, AdaptFormer can be defined as $\bm{h_{9}}=\bm{h_{9}}+\operatorname{Adapter}\bm{(h7)}$ where $\operatorname{Adapter}$ follows Equation 14.
ConvPass
(Convolutional By-Passes) [40] addresses the concern that many existing Adapters lack visual inductive bias, potentially limiting their performance for downstream vision tasks with limited data. To this end, the authors introduce a convolutional bottleneck module, running in parallel with the MSA or(and) MLP block. This module encompasses a $1\times 1$ convolution reducing the channel with $\bm{W}_{\text{down }}\in\mathbb{R}^{r\times D}$, a $3\times 3$ convolution with the same input and output channel, a $1\times 1$ convolution expanding the channel $\bm{W}_{\text{up}}\in\mathbb{R}^{D\times r}$, two nonlinear functions $\sigma$ and a scaling factor $s$, as shown in 9(b). The authors argue that Convpass is more efficient at capturing visual information in low-data scenarios due to its hard-coded locality of convolutional layers. The formal definition of Convpass is shown in Equation 15.
$\begin{gathered}h_{5}=\operatorname{Convpass_{1}}(h_{2})+h_{5}\quad h_{9}=\operatorname{Convpass_{2}}(h_{7})+h_{9}\\ \operatorname{Convpass}(h)=s\cdot\bm{W}_{\text{up}}\sigma(\operatorname{Conv2d}(\sigma(\bm{W}_{\text{down }}h)))\end{gathered}$ (15)
RepAdapter
[61] found that the removal of the nonlinear function in the Adapter does not result in performance degradation for vision tasks. In light of this finding, the authors propose a linear Adapter with group-wise transformation [62] and sequentially added two of these linear Adapters to both MSA and MLP blocks. Owing to the sequential placement of the RepAdapter and its inherent linearity, the additional parameters can be re-parameterized to the original MSA or MLP block after training, thereby incurring zero additional costs during inference. RepAdapter is illustrated in 9(c) and formally defined in Equation 16.
$\begin{gathered}h_{5}=\operatorname{RepAdapter_{1}}(h_{2})\quad h_{7}=\operatorname{RepAdapter_{2}}(h_{7})\\ \operatorname{RepAdapter}(h)=s\cdot\phi_{\text{up}}(\phi_{\text{down }}(h))+h\\ \tilde{h}=\phi_{\text{down }}(h)=\bm{W}_{\text{down }}h\\ \phi_{\text{up}}(\tilde{h})=[\bm{W}_{g1}\tilde{h}_{g1},\ldots,\bm{W}_{gG}\tilde{h}_{gG}]\\ \end{gathered}$ (16)
where $\bm{W}_{\text{down }}\in\mathbb{R}^{r\times D}$, $\tilde{h}_{g(1,\ldots,G)}\in\mathbb{R}^{\frac{r}{G}\times(N+1)}$ is the features splitted from $\tilde{h}\in\mathbb{R}^{r\times(N+1)}$ and $G$ is the number of groups in group-wise transformation [62]. $\bm{W}_{g(1,\ldots,G)}\in\mathbb{R}^{\frac{D}{G}\times\frac{r}{G}}$ is the projection weight matrix.
(a) Adapter
(b) Convpass
(c) RepAdapter
Figure 9: Comparison of three Adapter structures.
B.5 Selective Parameter Tuning Methods
The methods falling within this category aim to selectively update the parameters of a pre-trained model for downstream tasks. Within transfer learning, two prominent strategies, namely full fine-tuning and linear probing [45, 105], represent the two extremes of this category. Full fine-tuning updates all the model parameters end-to-end based on the new dataset while linear probing treats the pre-trained model as a feature extractor and only updates the prediction heads while keeping the backbone frozen. Although full fine-tuning generally exhibits superior performance compared to linear probing [99], it possesses certain limitations that may hinder its practicality in real-world production settings. Firstly, it requires running gradient descent for all parameters and necessitates storing a separate fine-tuned model for each task, incurring significant computational, memory, and storage overhead. These challenges become more salient with Transformer-based models whose parameters grow exponentially. Secondly, full fine-tuning may distort pre-trained features and underperform linear probing in out-of-distribution (OOD) scenarios [47].
To cope with the above issues, a cohort of PETL methods has emerged under this category. In addition to the two common approaches mentioned above, our investigation encompasses seven methods that can be further categorized into two groups: direct selective tuning [98, 5, 93], which involves the direct modification of selective weights, and efficient selective tuning [34, 41, 53], which approximates the weight updates with low-rank factors.
Notably, an extra advantage of methods in this category is that they introduce no additional inference latency, making them particularly favourable when inference efficiency is a priority. Methods within the direct selective tuning group abstain from introducing any new modules, thus inherently avoiding extra inference latency. Meanwhile, for methods in the efficient selective tuning group, the added modules can often be seamlessly integrated into weights of the pre-trained models through the re-parameterization techniques [37, 16], thereby ensuring the absence of increased inference latency as well.
B.5.1 Direct Selective Tuning
BitFit
[98] is a simple yet effective method that only tunes the bias parts of the pre-trained model. For each Transformer layer in ViT, BitFit updates the bias terms in the QKV projections and the FC layer in the MSA block, two FC layers in the MLP block and two LN blocks. It also updates the bias in the projection for patch embedding. The original authors underscore BitFit’s capability to achieve performance comparable to full fine-tuning or even surpass it under low and medium-data scenarios in BERT models [43].
LayerNorm
[5] represents another simple but strong baseline that solely tunes the two LN blocks in each Transformer layer - one before the MSA block and another before the MLP block. Given that each LN block contains merely two trainable parameters $\{\bm{W}_{LN},\bm{b}_{LN}\}\in\mathbb{R}^{D}$, LN-tune stands out as an exceedingly light-weight approach compared to other PETL methods. For instance, ViT-B/16 ($\sim$86M parameters) has only $\sim$38k LN parameters, accounting for $\sim$0.04of the total parameters.
DiffFit
[93] is a recently proposed PETL strategy designed for adapting large pre-trained diffusion models to the new domains. DiffFit exclusively fine-tunes the bias terms and the LN blocks within the network. Furthermore, it inserts learnable scale factors $\gamma$ to shift the features after the MSA and the MLP blocks, as shown in Equation 17. Conswasequently, DiffFit can be regarded as a combination of the BitFit and Ln-Tune, incorporating additional feature shift factors.
$\begin{gathered}{h_{5}}=\gamma_{1}\cdot{h_{5}}\\ {h_{9}}=\gamma_{2}\cdot{h_{9}}\end{gathered}$ (17)
B.5.2 Efficient Selective Tuning
LoRA
(Low-Rank Adaptation) [34] drew inspiration from recent investigations demonstrating that the learned over-parametrized models in fact reside on a low intrinsic dimension [50, 1]. Building upon this insight, the authors hypothesize that the change in weights during model adaptation also exhibits a low intrinsic rank and injects trainable low-rank decomposition matrices to approximate the weight updates. The LoRA update methodology is strategically applied to the Query/Value projection weights $\bm{W}_{Q/V}\in\mathbb{R}^{D\times D}$ within the MSA block. Concretely, the weight updates are approximated as $\bm{W}_{Q/V}+\Delta\bm{W}_{Q/V}=\bm{W}_{Q/V}+\bm{W}^{Q/V}_{\text{down}}\bm{W}^{Q/V}_{\text{up}}$ where $\bm{W}^{Q/V}_{\text{down}/\text{up}}\in\mathbb{R}^{D\times r/r\times D}$ and rank $r\ll D$. The authors use a random Gaussian initialization for $\bm{W}^{Q/V}_{\text{up}}$ and zero for $\bm{W}^{Q/V}_{\text{down}}$ so that $\Delta\bm{W}_{Q/V}=\bm{W}^{Q/V}_{\text{down}}\bm{W}^{Q/V}_{\text{up}}$ is zero at the beginning of training. The formal definition of LoRA is articulated in Equation 18, utilizing the notations delineated in Figure 8.
$\begin{gathered}{h_{3}}=\operatorname{LoRA}({h_{2}})+{h_{3}}\\ h_{3}=[\bm{Q},\bm{K},\bm{V}]\\ \operatorname{LoRA}({h_{2}})=[\bm{W}^{Q}_{\text{down}}\bm{W}^{Q}_{\text{up}}h_{2},0,\bm{W}^{V}_{\text{down}}\bm{W}^{V}_{\text{up}}h_{2}]\end{gathered}$ (18)
FacT
(Factor Tuning) [41] is inspired by the recent advances in Transformer compression [89, 101]and exploited the low-rank update paradigm (e.g., LoRA) to the extreme. While LoRA posits that the update for an individual weight matrix manifests a low-rank characteristic during fine-tuning, FacT advances the proposition that the weight updates spanning different matrices can also be effectively approximated using low-rank decomposition matrices. Specifically, FacT encapsulates the four weight matrices $\bm{W}_{Q/K/V/O}\in\mathbb{R}^{D\times D}$ in the MSA block and the two weight matrices $\bm{W}_{1}\in\mathbb{R}^{D\times 4D}$, $\bm{W}_{2}\in\mathbb{R}^{4D\times D}$ in the MLP block into a single $\bm{W}_{FacT}\in\mathbb{R}^{12M\times D\times D}$ tensor where $M$ is the number of Transformer layer. The update of $\bm{W}_{FacT}$, $\Delta\bm{W}_{FacT}$, can be decomposed into several factors to promote parameter efficiency. To this end, the authors leverage the well-established Tensor-Train (TT) [70]and the Tucker (TK) [12] format to decompose $\Delta\bm{W}_{FacT}$. $\text{FacT}_{\text{TT}}$ and $\text{FacT}_{\text{TK}}$ are used to denote different decomposition formats for FacT and their formal definitions can be found in Equation 19.
$\text{FacT}_{\text{TT}}:\Delta\bm{W}_{FacT}=s\cdot\bm{\Sigma}\times_{2}\bm{U}^{\top}\times_{3}\bm{V}^{\top}$ (19)
$\text{FacT}_{\text{TK}}:\Delta\bm{W}_{FacT}=s\cdot\bm{A}\times_{1}\bm{B}^{\bm{\top}}\times_{2}\bm{U}^{\bm{\top}}\times_{3}\bm{V}^{\bm{\top}}$ (20)
where $\bm{U}\in\mathbb{R}^{D\times r},\bm{V}\in\mathbb{R}^{D\times r},\bm{\Sigma}\in\mathbb{R}^{12L\times r\times r},\bm{B}\in\mathbb{R}^{12L\times r}\bm{A}\in\mathbb{R}^{r\times r\times r}$ and the ×$\times_{j}$ denotes mode-$j$ product and $s$ is the scaling factor.
Since $\Delta\bm{W}_{FacT}$ contains the updates for $\bm{W}_{Q/K/V/O},\bm{W}_{1/2}$, the modified forward pass inherently influences $h_{3},h_{5},h_{8},h_{9}$. Let’s consider $h_{5}$ for elucidation. Once the weight update $\Delta\bm{W}_{FacT}$ is calculated with $\text{FacT}_{\text{TT(TK)}}$ in Equation 19, the corresponding update for $\bm{W}_{O}$, $\Delta\bm{W}_{O}$, is extracted from $\Delta\bm{W}_{FacT}$. Similar to the modified forward pass of LoRA, $h_{5}=h_{4}\Delta\bm{W}_{O}+h_{5}$.
SSF
(Scale & Shift deep Features) [53] employs linear transformations to adapt the intermediate features extracted by a pre-trained model. Motivated by the feature modulation methods [36, 71], SSF is designed to accommodate the distribution difference between the upstream and downstream datasets. Specifically, SSF modulates the features residing at $h_{2},h_{3},h_{5},h_{7},h_{8},h_{9}$ by incorporating scale and shift factors. To demonstrate the mechanism of SSF, let’s consider $h_{5}\in\mathbb{R}^{(N+1)\times D}$ as an illustrative example and other features can similarly undergo the same transformative process. Formally, the modulated $h_{5}$ is formulated as follows.
$\begin{gathered}h_{5}=\operatorname{SSF_{5}}(h_{5})=\bm{w}^{5}\odot h_{5}+\bm{b}^{5}\end{gathered}$ (21)
where $\bm{w}^{5}\in\mathbb{R}^{D}$, $\bm{b}^{5}\in\mathbb{R}^{D}$ are the scale and shift factors affiliated with the SSF module attributed to $h_{5}$, and $\odot$ is the dot product. It is noteworthy that each modulated feature has its own SSF module with corresponding scale and shift factors. The modification details for other features are summarized in Table 5.
Appendix C More Results
Drop-path-rate.
Learning with low-shot data is prone to over-fitting. We find that if the drop path rate — which stochastically drops a transformer block per sample [35] — is set not as default (i.e., nonzero), all the methods can benefit from such a regularization. Figure 10 shows the performance gain by tuning the drop-path-rate on compared with the default 0.
Figure 10: Performance gain for PETL methods by turning drop-path-rate on.
More results on prediction similarity analysis.
We provide prediction analysis discussed in section 4 for all the datasets in VTAB-1K. It is expected that their predictions are similar for datasets with very high accuracy, such as Flowers102 (avg 99.1%) and Caltech101 (avg 91.4%). Beyond them, we find that most PETL methods show diverse predictions in other datasets in VTAB-1K.
(a) Caltech101
(b) Camelyon
(c) Cifar100
(d) Clevr-Count
(e) Clevr-Dist
(f) DMLab
(g) dSpr-Loc
(h) dSpr-Ori
(i) dSpr-Ori
(j) dSpr-Ori
(k) DTD
(l) EuroSAT
(m) Flower102
(n) KITTI
(o) Pets
(p) Resisc45
(q) Retinopathy
(r) sNORB-Azim
(s) sNORB-Elev
(t) Sun397
(u) SVHN
Figure 11: Prediction similarity analysis on other datasets.
WiSE PETL results for all distribution shift datasets.
We provide detailed WiSE PETL performance for each distribution shift dataset in Figure 12. WiSE improves both the robustness and the in-distribution performance of PETL methods. Interestingly, even though full fine-tuning is generally less robust than PETL methods, applying WiSE allows it to achieve better performance in both in-distribution and OOD data.
(a) ImageNet-A
(b) ImageNet-R
(c) ImageNet-S
(d) ImageNet-V2
Figure 12: WiSE PETL performance on all distribution shift datasets. In-distribution (x-axis) vs. OOD (y-axis)
Appendix D Broader Impacts
Our study provides a unifying study of PETL in visual recognition. We expect it to serve as a valuable practical user guide to benefit society. Specifically, fine-tuning large models needs significant computation. A unifying study of PETL will ease end-users to apply more parameter-efficient and computation-efficient ways for fine-tuning. To our knowledge, our paper does not introduce any additional negative societal impacts compared to existing papers on PETL.